The rate constant of a reaction depends on


a)      Temperature


b)      Initial concentration of the reactants


c)       Time of reaction


d)      Extent of reaction


Grade:11

7 Answers

Navjyot Kalra
10 years ago

a)      Temperature

Exp:       It is a constant of particular reaction at a given temperature. It does not depend upon initial concentration of the reactants time of reaction and extent of reaction.

meenakshi boni
10 years ago

A

hanuman balagoni
10 years ago

the rate constant depends on initial conc.of the reactants because by the formula of rate constant=initial conc./time taken to convert the reactants into products. It  depends on time also.

ANIL
10 years ago

a.temperature

Nandakumar UK
10 years ago

The rate constant will change with Temperature, use of catalyst and also when the unit of rate constant is changed. For zero order reaction ty/x =(y/x) C0/k or k = (y/x) C0/ty/x. Here also k does not change with change in C value. 

ashu
10 years ago
Rate constant is directly proportional to temperature because as temp increases velocity of molecule increases and rate contant also increases
EPARI SRAGVEE
10 years ago
answer is `a` rate constant depends on temp. but not anuy other factor such as intial concentration.according to the formula it is the concentration at equilibrium